### v4

# Unforgettable Lessons from Forgettable Images: Intra-Class Memorability Matters in Computer Vision

Abstract
We introduce intra-class memorability, where certain images within the same class are more memorable than others despite shared category characteristics. To investigate what features make one object instance more memorable than others, we design and conduct human behavior experiments, where participants are shown a series of images, and they must identify when the current image matches the image presented a few steps back in the sequence. To quantify memorability, we propose the Intra-Class Memorability score (ICMscore), a novel metric that incorporates the temporal intervals between repeated image presentations into its calculation. Furthermore, we curate the Intra-Class Memorability Dataset (ICMD), comprising over 5,000 images across ten object classes with their ICMscores derived from 2,000 participants’ responses. Subsequently, we demonstrate the usefulness of ICMD by training AI models on this dataset for various downstream tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing. Surprisingly, high-ICMscore images impair AI performance in image recognition and continual learning tasks, while low-ICMscore images improve outcomes in these tasks. Additionally, we fine-tune a state-of-the-art image diffusion model on ICMD image pairs with and without masked semantic objects. The diffusion model can successfully manipulate image elements to enhance or reduce memorability. Our contributions open new pathways in understanding intra-class memorability by scrutinizing fine-grained visual features behind the most and least memorable images and laying the groundwork for real-world applications in computer vision. We will release all code, data, and models publicly.
1 Introduction
Figure 1: An object instance can be more memorable than another within the same class. Intra-class memorability refers to the ability of instances within the same category to be remembered. First, we present example images for the five exemplar classes. Images are identified as low (Row 1) or high (Row 2) memorability based on results from human behavior experiments. We then use these high- and low-memorability images to train and test state-of-the-art AI models across four downstream computer vision tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing.
As we experience the world, our memory selectively retains certain objects while overlooking others [1, 2, 3]. For instance, among numerous flowers we encounter, a pale white daisy might fade from memory, while a striking violet osteospermum with intricate shapes remains vividly etched (Fig. 1). What factors make some flowers more memorable than others? To answer this question, we first introduce intra-class memorability which refers to the ability of instances within the same object class to be remembered. To delve into why certain items are more memorable than others, despite belonging to the same overall category, we conduct human behavior experiments using a continuous recognition task. In this task, participants are presented with a sequence of new and repeated images from the same class with delayed repetition and they are asked to identify those repeats. The proposed paradigm is novel compared to existing research, where continuous recognition tasks typically involve sequential presentations of images from different classes [4, 5, 6, 7, 8, 9].
To measure the memorability of individual instances, we introduce the Intra-Class Memorability score (ICMscore). Unlike prior evaluation metrics [7, 8, 10, 11], our ICMscore incorporates average hit and miss rates for an item’s repetitions, weighted by the time intervals between its initial and repeated presentations. Missing a repeat after a short interval indicates low memorability (LM), while remembering a repeat after a long interval signifies high memorability (HM).
Due to the lack of intra-class memorability datasets, we contribute the Intra-Class Memorability dataset (ICMD). The dataset contains 10 classes spanning over 500 images per class with their corresponding ICMscore aggregated from the responses of 2000 participants in the human behavior experiments on continuous recognition tasks. To demonstrate the effectiveness of our ICMD, we train and test AI models on different subsets of ICMD images based on their ICMscores across four downstream computer vision tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing. Our experimental results show that a memorability prediction model trained on ICMD captures representative features diagnostic for intra-class memorability. Moreover, surprisingly, while highly memorable images seem more salient and thus easier to remember, we found that these images actually impair AI performance in image recognition and continual learning tasks. Finally, we use images from ICMD with and without masked semantic objects to fine-tune the state-of-the-art image diffusion model [12] using LoRA [13] to control image memorability. We highlight our key contributions:
1. We introduce the novel and important problem of intra-class memorability and provide insights into why certain instances are more memorable than others, despite belonging to the same category.
2. We challenge existing paradigms for collecting memorability data on naturalistic images by introducing a new human behavior experiment on intra-class memorability. In addition, we propose a novel memorability metric, the ICMscore, which accounts for the time intervals before a repeated image is identified in continuous image recognition.
3. We collect ICMscore from the responses of 2000 participants on 5000 images spanning across 10 object classes and contribute the first dataset on intra-class memorability, named as Intra-Class Memorability dataset (ICMD). To demonstrate the usefulness of our dataset, we introduce four computer vision tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing.
4. Our results show that our memorability prediction model is capable of capturing key features for intra-class memorability. Surprisingly, while HM images are more salient and tend to be retained longer in memory, training AI models on these images can actually impair their performance in image recognition and continual learning tasks. Finally, trained on ICMD images with and without their semantic objects masked out, our fine-tuned state-of-the-art image diffusion model can enhance or reduce the memorability of a given image by manipulating their semantic regions.
2 Related Work
Research in cognitive psychology indicates that certain visual and contextual features—such as prominent visual characteristics, emotional valence, and semantic richness—substantially increase the likelihood of remembering specific objects over others [14]. This implies that memorability is a predictable and universal property, rather than purely subjective [15, 16, 17]. Based on this, Isola et al. [17] define “memorability" as the probability that a viewer will remember an image after a single viewing. Subsequent studies have investigated how intrinsic and extrinsic factors influence memorability. Extrinsic factors refer to an image’s categorical information, while intrinsic factors include other visual features beyond these categorical information [5, 6, 9]. However, none of these studies have effectively disentangled the effects of extrinsic and intrinsic features on memorability.
To address this problem, subsequent works [4, 7, 8, 18, 19] focus on specific super-categories, such as animals or indoor and outdoor scenes. In addition, some research has extended the studies on static images to videos, exploring memorability differences across video super-categories [20, 21, 22]. Although these studies limit categories to reduce the impact of extrinsic factors on memorability, they still mix images from different categories, making it difficult to conduct controlled studies that solely examine the effect of intrinsic features on memorability. To address this research gap, we introduce the concept of intra-class memorability. To investigate it, we propose a novel human behavior experimental paradigm focusing memorability among images within the same class.
Several memorability datasets have been collected through human behavior experiments [17, 4, 6, 7, 8], with LaMem [5] as one of the largest for deep learning-based image memorability prediction. In these studies, participants typically view a sequence of images, identifying any repeated target images. Memorability scores are often calculated based on the proportion of participants who correctly identify these repeats [17, 4, 6]. Some studies adjust scores to account for false alarms, where images are mistakenly identified as repeats [23, 7, 8]. The time interval between target presentations also influences memorability [5], yet random sampling of intervals can lead to unfair comparisons due to inconsistent controls. In our experiment, we standardize time intervals and introduce a novel memorability score, named as Intra-Class Memorability score (ICMscore). It accounts for both hit rates and miss rates of repeated images across a predefined set of time intervals. Moreover, we contribute a new dataset, Intra-Class Memorability dataset (ICMD). It consists of 5000 images with their corresponding ICMscores based on responses from over 2000 participants.
Several studies have demonstrated that HM images can positively influence human cognitive states. For example, GANalyze [24] introduces a generative model that modifies image attributes to affect human emotional and cognitive responses. Peng et al. [25] show that memorable images boost user engagement in image retrieval systems and improve recall accuracy in storytelling and educational contexts. Similarly, [26] and [27] explore how AI-modified images and videos uniquely shape human memory. Thus, predicting memorability for a given image has become increasingly important [28, 29, 30]. However, existing models for predicting memorability rely on both extrinsic and intrinsic features, making them unable to capture fine-grained memorability differences within the same class of images. In contrast, our experiments demonstrate that our model, trained on the ICMD, outperforms existing predictors in accurately predicting intra-class memorability.
Despite early efforts to explore the role of memorability in enhancing human cognition, its impact on computer vision tasks remains largely unexplored. Specifically, the connection between image features relevant to object recognition and those that explain memorability is still unclear. Understanding this relationship could have significant implications for fields like continual learning. Using our ICMD, we present preliminary evidence that memorability affects AI performance in tasks such as image recognition and continual learning. To edit images with controls on memorability, we also build upon the existing generative models [12, 13] and introduce an image diffusion model, capable of manipulating image elements to enhance or reduce memorability.
3 Intra-Class Memorability Dataset(ICMD)
Figure 2: Schematic illustration of the human behavior experiment on intra-class memorability. A sequence of new and repeated images is presented. Each image appears for 1200 ms, followed by a blank screen for 1600 ms. Participants are asked to identify those repeats by pressing the space on the keyboard. Red arrows indicate the correct identification of the repeated images. Unique foil images are inserted into the sequence whenever there is no target image in that location.
3.1 Human Behavior Experiments
Different from the existing memorability datasets [4, 5, 6, 9, 7, 8], our work introduces a new Intra-Class Memorability Dataset (ICMD), comprising naturalistic images from ten distinct object categories along with corresponding human memorability responses. The selected categories—volcano, dog, phone, lemon, car, bird, palace, flower, jellyfish, and teapot—were chosen to ensure clear semantic separation with no category overlap. We curated a total of 5,000 target images from ImageNet [31], evenly distributed across these ten classes. Fig. 2 presents the schematic of the human psychophysics experiments in the continuous object recognition tasks. In each trial, participants view a sequence of images, with both target and foil images belonging to the same category. Each image is followed by a blank screen to allow ample response time. Participants press the space bar when recognizing a previously seen image, with reaction times recorded during the 1,200 ms image presentation and the subsequent 1,600 ms blank screen. Before the start of an experiment, participants complete a guided demonstration to ensure task comprehension. During the experiment, immediate feedback follows each response: a green tick for correct answers and a red cross for incorrect ones. The image sequence is randomly shuffled, with target images reappearing after intervals of $t\in[8,16,24,32]$ images, uniformly distributed across these intervals. We determine these interval ranges through a pilot study exploring intervals $t$ from 2 to 64. The results reveal that intervals shorter than 8 led to near-ceiling performance, with participants achieving over 93% recall accuracy over all target images. Conversely, intervals exceeding 32 proved too challenging, with more than 34% of participants failing to identify any repeated images. Based on these findings, we select the interval range of [8, 16, 24, 32] to ensure meaningful image-dependent variability in memory recall performance. Each trial includes 100 unique target images, with foil images filling the sequence when targets are not shown, each target image and its repeat are shown only once to one participant. Sec. A.1 and Fig. A1 highlight differences in the experiment schematic between our study and previous works.
A total of 3,263 participants are recruited on Amazon Mechanical Turk (MTurk) [32]. To ensure data quality, participants are disqualified if they incorrectly identify more than three foil images as target images in a row or fail to respond within 60 seconds. After filtering the data from disqualified participants, we analyze responses from the remaining 2,000 qualified participants, resulting in 4,000 trials and a total of 346,537 key presses, with an average of 85 key presses per trial. The average false positive rate of the key presses is 15.24%. All experiments are conducted with informed consent and follow protocols approved by our institution’s Institutional Review Board.
3.2 Intra-Class Memorability Score (ICMscore)
To assess intra-class memorability, we introduce the Intra-Class Memorability Score (ICMscore), which measures the memorability of an image independently of its categorical information. Previous research typically computes memorability scores using hit rates and miss rates of memory recall [23, 7, 8]. However, the interval $t$ between image repetitions significantly affects memorability [5]. As shown in Fig. A2 and Sec. A.2, memory accuracy declines with longer intervals due to memory decays and increasing cognitive loads over time [33, 34]. Therefore, we incorporate the lengths of these time intervals into the ICMscore calculation.
For each target $I$, we receive key press responses from $n$ participants. The correctness of $i$th participant’s response is denoted as $x_{i}$, forming a response set $\{x_{1},x_{2},\dots,x_{n}\}$ for image $I$. A response of $x_{i}=1$ indicates a correct identification, while $x_{i}=-1$ indicates an incorrect identification. The interval associated with each response $x_{i}$ is represented as $t_{i}$.
Since human memory retention diminishes over longer intervals, a longer interval $t$ makes correct identification of seen images more challenging. Therefore, when a participant correctly identifies an image after a longer interval, we assign it a higher memorability score $s$ to reflect the increased difficulty associated with the larger $t$. For correct responses ($x_{i}=1$), the memorability score $s_{i}$ is computed using a linearly increasing function: $s_{i}=x_{i}\frac{t_{i}}{T}$, where $t_{i}$ is the interval between two presentations of a target image, and $T=32$ is the maximum interval. Conversely, as longer time intervals are more prone to incorrect identifications of repeated target images, errors during shorter intervals are more indicative of the image’s memorability. For incorrect responses ($x_{i}=-1$), shorter intervals incur a larger penalty, while longer intervals result in a smaller penalty. The memorability score for incorrect responses is given by $s_{i}=x_{i}\frac{T-t_{i}}{T}$. The final ICMscore for each target image $I$ is calculated as the average of its memorability scores $s_{i}$ over all $n$ participants: ICMscore(I)$=\frac{\sum_{i=1}^{n}s_{i}}{n}$.
To highlight the importance of the temporal penalty in our ICMscore formulation, we conducted an additional analysis by removing the temporal component $t_{i}$. See Sec. A.5 for details. Results show that without this penalty, ICMscore fails to capture intra-class memorability effectively, diminishing its utility for downstream computer vision tasks. Furthermore, to demonstrate that ICMscore is not merely a variant of existing memorability metrics defined in [23, 7, 8], we applied an existing memorability prediction model trained on LaMem dataset [5] to estimate memorability scores for images in our ICMD dataset. The two sets of scores show only a weak positive correlation, with a Pearson coefficient of 0.26. This suggests that ICMscore captures distinct memorability-related features overlooked by prior metrics and serves a complementary role in memorability evaluation.
Figure 3: Correlation analysis and Grad-CAM visualization. (a) Pearson correlation between feature distances of images to their categorical centroids and their ICMscore. The y-axis shows correlation values while the x-axis displays class labels. Correlations range from $r=0.21$ for “Flower" to $r=0.21$ for “Phone" ($p<0.05$ for all ten classes). (b) Grad-CAM visualizations displaying importance maps for three high memorability (HM) images. These visualizations highlight key regions by computing gradients of the memorability predictor with respect to feature maps to identify areas contributing most to memorability scores. For HM images, features enhancing memorability are highlighted according to the color scale shown in the right color bar. The first column presents original images, the second column shows visualizations from the model trained on LaMem dataset [5], and the third column displays results from our model trained on ICMD.
3.3 Dataset Analysis
Given the target images and their corresponding ICMscores, we now analyze how visual features of target images influence ICMscores. To investigate this relationship, we use a ResNet [35] model pre-trained on ImageNet [31] for object recognition tasks to extract feature representations for each image. We define the mean of all image features belonging to the same category as the centroid. Next, we compute the Pearson correlation [36] between the L2 distances from each centroid to individual target image features and their ICMscores.
Our results reveal a positive correlation between an image’s distance from its categorical centroid and its ICMscore across all object classes, as shown in Fig. 3(a). The further away the features of an image are from its categorical centroid, the more memorable the image becomes. Here, we explore the connections between memorability and feature distance within categories. Insights gained from this intra-class feature analysis motivate us to design the experiments that investigate the role of memorability in the following computer vision tasks in Sec. 4.
4 Experiments
4.1 Memorability Prediction
Memorability prediction can provide insights into human cognition and perception by helping us understand how and why certain images are more memorable than others [37, 38, 39]. Existing memorability prediction methods fail to disentangle the intrinsic and extrinsic factors that influence the memorability of an image. In contrast, our work focuses on predicting ICMscore based exclusively on intrinsic factors of images within the same category.
Specifically, we use a ResNet [35], pre-trained on ImageNet [31] for image recognition tasks, to extract image features, followed by two fully connected layers to predict memorability scores. We train our ICMscore predictor using the ICMD with an L2 regression loss, where the ICMscore of the input image serves as the ground truth. To validate the effectiveness of our predictor, we compare it against two baseline models. Baseline 1 employs the same network structure as our model but is trained from scratch without loading the pre-trained network parameters. For Baseline 2, inspired by previous work [30, 40], we use a Vision Transformer (ViT) pretrained on ImageNet [41], append two randomly initialized fully connected layers, and fine-tune only these added layers while keeping the ViT backbone frozen. We perform 5-fold cross-validation to ensure the statistical robustness of the results. Following the practices in memorability prediction [10, 30, 5], we assess the models’ performance by measuring the Spearman rank correlation [42] between the predicted and ground truth ICMscores. The first model achieves the highest correlation (0.64 ± 0.09), demonstrating that ICMscores can be effectively predicted using features extracted from pre-trained image recognition models, while the other two models obtained lower correlations: 0.60±0.2 and 0.57±0.05.
To ensure the 0.64 correlation from our predictor is not due to model overfitting, we split human participants into two independent groups and computed the Spearman correlation of their ICMscores on the same target images. Across 10 runs, the average correlation is $0.69\pm 0.05$, demonstrating high consistency in within-human memorability. Although our model’s correlation is slightly lower than human consistency, it exceeds both baselines, suggesting that the model effectively captures diagnostic visual features of memorability rather than overfitting the data. This finding also aligns with the observed positive correlation between feature distance from the categorical centroids and ICMscores (Sec 3.3).
To further explore what image components contribute to high memorability, we employ Grad-CAM [43] to visualize importance maps for memorability, as shown in Fig. 3(b). For highly memorable (HM) images, our predictor focuses on distinctive regions. For instance, in a teapot image with an ICMscore of 0.78, our predictor emphasizes the unique face-like texture of the teapot, while the model trained on LaMem [5] focuses on the teapot’s spout and handle which are typical features reflecting the generic categorical appearance of teapots.
4.2 Image Recognition
Figure 4: Performance of AI models using HM, MM, and LM images for image recognition task (a) and continual learning with replay (b) and (c). (a) Performance comparison in Top-1 Accuracy between image classification models trained on HM (green), MM (yellow), and LM (blue) images is presented. The chance level is 10% for the 10-way classification (red dashed line). Error bars represent standard errors calculated over 5 independent runs. One-way ANOVA across all models yields $p<0.05$, indicating statistically significant accuracy differences among green, yellow, and blue bars. (b) Performance of continual learning models using HM (green), MM (yellow), and LM (blue) images for replays, is reported as Accuracy (left), BWT (middle), and Forgetting (right). See Sec. 4.3 for details on evaluation metrics and continual learning models. Error bars represent standard errors calculated over 5 independent runs. Two-tailed t-test analysis comparing HM vs. LM models and MM vs. LM models across all three metrics shows significant differences ($p<0.05$). (c) Accuracy at the last task (solid lines) and Forgetting (dotted lines) as a function of task numbers for continual learning models using HM (green), MM (yellow), and LM (blue) images for replays. The y-axis (Score) represents either Accuracy or Forgetting as indicated in the legend.
We explore the role of memorability in image recognition by training models on image subsets with varying ICMscores. Images within each object class are sorted by their ICMscores and divided into two subsets: HM (High Memorability) with the top 250 images and LM (Low Memorability) with the bottom 250. Besides, we create a Mixed-Memorability (MM) subset of 250 images randomly mixing HM and LM images from the ICMD.
We employ ResNet [35] as the backbone with random weights, and train it from scratch on each of the three curated subsets above over all ten classes using the Adam optimizer [44]. The initial learning rate is set to $1\times 10^{-3}$ and reduced by 50% every five epochs over a total of 20 epochs following a linear decay schedule. The resulting models are labeled as HM-model, MM-model, and LM-model based on their training subsets.
We use the Cross-Entropy loss to train all three models for the 10-way classification task. Test images for object recognition are obtained from the original ImageNet [31] test set regardless of their ICMscores. As these test images lack ground truth ICMscores, we could not evaluate the top-1 accuracy of the models on individual subsets of test images stratified by ICMscores. Fig. 4(a) presents the average top-1 accuracies of the three models on all the test images for all 10 object classes over 5 random seeds.
Our results in Fig. 4(a) indicate that the LM-model consistently outperforms the HM-model, achieving an average accuracy of $0.80$ compared to $0.76$ across all 10 object classes. Statistical analysis using one-way ANOVA reveal significant differences in top-1 accuracy between the two models ($p<0.05$). Surprisingly, the LM-model also surpasses the MM-model. These findings indicate that image memorability impacts image recognition, with LM images improving accuracy and HM images impairing it. We also provide the confusion matrices for the classification performances of all three models in Sec. A.3.
In real-world applications, training image recognition models in computer vision benefits from using images with varying memorability. However, obtaining memorability scores for all images through human behavioral studies is often time-consuming, costly, and therefore impractical. Therefore, a computational model for predicting an image’s memorability becomes invaluable. We leverage the ICMscore predictor, introduced in Sec. 4.1, to estimate the ICMscore for all images and then use these predictions as criteria to conduct the same experiment for image recognition. From Fig. A5 (a), consistent with the findings above, recognition models trained exclusively on LM images categorized by predicted ICMscores achieve higher top-1 accuracies than the MM-model. One possible reason is that easy images for object recognition often feature prototypical representations of object classes [45], which accelerate training and foster robust, generalizable learning. As shown in Sec 3.3, easy images tend to be nearer to categorical centroids, and hence, less memorable but more beneficial for representation learning in AI models.
For comparison, we also conduct the same image recognition experiment based on the classical memorability definitions from prior works [4, 5, 6, 7, 8] to divide the dataset into two subsets and utilize the predictor from [5]. The results, shown in Fig. A5 (b), indicate that the predictor proposed by [5] does not enhance image recognition accuracy. Additionally, we observe no significant difference in top-1 accuracy between models trained on the HM and LM images as predicted by [5].
4.3 Continual Learning
Continual learning, particularly class-incremental learning, is essential for AI models to adapt to new classes over time without experiencing catastrophic forgetting of previously learned classes [46]. One type of continual learning methods is based on memory replay [47, 48, 45], which involves storing a subset of previously encountered examples in a memory buffer and replaying them during subsequent tasks. Given that some images are easier to remember while others are not, we conjecture that image memorability may play a key role in improving the continual learning performance of AI models. Here, we investigate how memorability can be leveraged to select exemplar images for replay, potentially enhancing the model’s ability to retain knowledge across tasks.
We structure our ICMD for continual learning by training object recognition models on the first 5 object classes in the initial task, followed by introducing 1 incremental class in each subsequent task. Since there are a total of 10 object classes, this results in 6 tasks. To examine the role of image memorability, we further divide the images from each class into HM, MM, and LM image subsets by following the same procedures in Sec. 4.2. During replay, we exclusively use images from one of these subsets, meanwhile keeping training images identical for all three continual learning models across all 6 tasks. We label the continual learning models with replays from HM, MM, and LM images as HM-models, MM-models, and LM-models respectively. We train all these 3 continual learning models in the initial task for 10 epochs, and incremental tasks for 5 epochs with an Adam optimizer using a learning rate of $1\times 10^{-3}$. The replay buffer size is set to 20 samples.
To benchmark the continual learning performance of these three models, we include the following three standard evaluation metrics, namely Accuracy at the last task (Acc), Backward knowledge Transfer (BWT), and Forgetfulness (F). We define $A_{t,i}$ as the accuracy of the model trained on all the tasks up to $t$ and tested on the test set of task $i$. Next, we introduce these metrics in details.
Accuracy at the last task (Acc) [49] is defined as the accuracy of the model trained up to the last task $T$ and tested on all the previous tasks. The higher the Acc, the better the model performs in continual learning tasks.
Backward Transfer (BWT)[50, 51, 49] quantifies the model’s knowledge retention ability by measuring performance drop on earlier tasks after learning new ones:
$BWT=\frac{\sum_{i=2}^{T}\sum_{j=1}^{i-1}\left(A_{i,j}-A_{j,j}\right)}{\frac{T(T-1)}{2}}$ (1)
Typically, the BWT is a negative value due to catastrophic forgetting of the model. The lower BWT, the worse the model performs in continual learning tasks.
Forgetting (F) [49] captures the accuracy drop on the test set of the first task for the model trained at the final task. It is defined as $F=A_{1,1}-A_{T,1}$. The lower the F, the better the model performs in continual learning tasks.
Results in Fig. 4(b) show that models using LM images for replay consistently outperform those using HM and MM images in Acc, BWT and F. We also plot accuracy and forgetting as functions of task numbers for all three models in Fig. 4(c). The results indicate that the model replayed with LM images consistently exhibits reduced forgetting and achieves higher accuracy across all tasks. Together with our feature analysis in Sec. 3.3, our continual learning results align well with the previous research [52, 53] that replaying samples where their features are closer to the categorical centroids yields better continual learning performance.
Similar to the image recognition experiments in Sec. 4.2, we categorize images into HM, MM, and LM using both our ICMscore predictor and the predictor from [5]. From Fig. A5, there is no significant performance difference in replay methods using LM and HM images predicted by [5]. In contrast, our ICMscore predictor remarkably reveals significant differences, with continual learning models replayed with LM images outperforming those with HM images or models using replay images predicted by [5].
4.4 Memorability-controlled Image Editing
Figure 5: Proof-of-conept demonstration of memorability-controlled image editing. (a) During training, we fine-tune the image diffusion model, Ip2p  [12], using LoRA [13] with pairs of masked images and their corresponding ground truth images. (b) During inference, SAM [54] segments and masks the teapot area of the source image, and our fine-tuned model takes the masked image as input to edit teapot images with varying memorability. (c) The high-memorability image generator increases the ICMscore of low-memorability (LM) teapots by adding striking textures and enhancing brightness. Conversely, the LM image generator reduces the ICMscore by blurring distinctive features, making the teapots appear more ordinary with plain textures. The numbers beneath the images represent ICMscores predicted by our memorability predictor as introduced in Sec. 4.1.
Editing images with controlled memorability offers valuable applications across various research fields, such as designing advertisements for commercial products and improving students’ knowledge retention ability in education [24, 55, 56]. Next, we present a proof-of-concept demonstration of using a subset of teapot images in our ICMD to train a state-of-the-art image diffusion model capable of modifying the memorability of an original teapot image.
To adjust a teapot image’s memorability while preserving its semantic content, we train a high-memorability teapot generator (HM generator) using a subset of teapot images with high memorability (HM images) calculated from human behavior experiments (Fig. 5a). First, for each HM image, we use SAM [54] to segment and mask the teapot area with white pixels. Next, these masked images are paired with their corresponding original images to train image diffusion models. Specifically, we fine-tuned the pre-trained InstructPix2Pix model (Ip2p) [12] with LoRA [13]. The generated images by Ip2p are compared against the ground truth HM images to improve its image generation quality.
During inference, the HM generator uses “highly memorable teapot" as the text prompt and the masked teapot image as input for image editing (Fig. 5c). The masked area guides the HM generator to focus on the specified region. The generator produces highly memorable teapot images. We qualitatively evaluate the HM generator’s performance using LM teapot images from our ICMD. As illustrated in Fig. 5c, the generator effectively enhances memorability by adding visually striking textures to the original teapots. For quantitative evaluation, we use our memorability prediction model (detailed in Section 4.1) to predict ICMscores for both the original and generated images. The results indicate that the generated teapots consistently achieve higher ICMscores than the originals, demonstrating the effectiveness of our HM generator for memorability-controlled image editing.
We also employ a similar training strategy to create a Low-Memorability teapot generator (LM generator), with the results shown in Fig. 5. Similar analyses as applied to the HM generator are conducted. The findings reveal that memorability increases with the addition of distinctive visual elements, while it decreases when the teapot’s unique textures are removed.
5 Discussion
We introduce intra-class memorability, defined as the ability of individual instances within the same image class to be remembered by humans. To investigate why certain instances are more memorable than others, we contributed the ICMD dataset, comprising 5,000 images across 10 object classes. Memorability for these images is quantified using the ICMscore metric, derived from human behavior experiments in continuous object recognition tasks. To showcase the utility of ICMD, we train and evaluate AI models on subsets of the dataset, categorized by ICMscore, across four downstream computer vision tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing. Our results reveal that intra-class memorability significantly impacts feature representation learning, with low-memorability images notably improving performance in image recognition and continual learning tasks.
Building on our promising results, we outline several key directions for future research. While our study demonstrates memorability’s role in computer vision using ten object classes, expanding ICMD to hundreds of ImageNet classes would provide broader insights. Furthermore, controlling image properties—such as standardizing backgrounds—could remove confounding factors and help isolate class-specific features critical for memorability. Although our findings highlight memorability’s impact on computer vision tasks, its underlying mechanisms remain an open question. Lastly, while we focus on supervised training, future work could explore intra-class memorability in self-supervised learning frameworks.
Appendix A Appendix
A.1 Difference in Human Behavior Experiments from the Existing Literature
Previous studies typically present participants with sequences of images where each sequence contains images from multiple categories. However, image category can significantly influence memorability [4, 7, 8, 18, 19]. For example, dog images are generally more memorable than mountain images [8]. To eliminate the influence of image categories on memorability, we design a new human behavior experiment focusing on intra-class memorability. Fig. A1 compares our experimental protocol with previous approaches [4, 5, 6, 9, 7]. Unlike previous methods that mix multiple categories in a single sequence, we separate image sequences by category. Each participant views image sequences from a single category, with breaks in between image sequences from different categories to reduce category interference on memorability.
Figure A1: The differences between our experimental protocol and previous studies. Our experimental design differs from previous studies [4, 5, 6, 9, 7]. In Row 1, “Others" represents the protocol used in previous studies, where participants are exposed to images from multiple categories within a single sequence. In Row 2, “Ours" represents our protocol, which requires participants to view only images from a single category per sequence, with the dotted line indicating the break between sequences. Different colored squares represent images from distinct categories, and the numbers indicate the position of each image within its category (e.g., B2 is the second image in category B). The black "X" represents a foil image that appears only once. In our protocol, the foil image "X" belongs to the same category as the target images in the same sequence.
A.2 The Effect of Time Interval in Human Behavior Experiments
In Sec. 3.2, we state that human memory retention tends to weaken over longer intervals. That is, a longer interval $t$ makes correct identification of the repeated images more challenging. See Sec. 3.1 for the definition of interval $t$. We demonstrate this effect in Fig. A2. Fig. A2 shows that identification accuracy of the repeated images and F1 score decrease with the longer interval $t$. Additionally, we conducted an experiment with $t$ exceeding 32, during which most participants struggled to complete the task. Therefore, we empirically set $t\in\{8,16,24,32\}$ in the actual experiments.
Figure A2: Human behavior experiment scores. This figure presents human performance at different intervals of $t$ (see Sec. 3.2 for the definition of $t$). The x-axis represents $t$, and the y-axis represents the score, with the red line indicating the accuracy score and the green line indicating the F1 score [57]. The values represent the average accuracy and F1 scores [57], computed across all trials with the same $t$. The error bars represent the standard error.
A.3 Confusion Matrices in Image Recognition Experiments
Following the image recognition experiment described in Sec. 4.2, we observe that models trained on LM images perform better in image recognition tasks than HM and MM images. To further demonstrate this difference, we compare the confusion matrices of models trained on three subsets. As shown in Fig. A3, the model trained on the LM images has a lower error rate across most classes, with the largest improvement in the “bird" class, where its accuracy increases by 16% and 15% in comparison to the HM and MM models.
Figure A3: Accuracy confusion matrices for models trained on HM, MM, and LM images classified based on human performances in behavior experiments. Each confusion matrix shows the proportion of classifying images as predicted labels (columns) given the ground truth (rows). Values in all the entries of the matrices represent the average classification proportion over 5 runs. See the color bar on the right for the classification proportion.
A.4 Performance of Models Using Memorability Predictors from the Previous Works
We replace ICMscore in ICMD with the memorability score from prior work [5] and split the dataset into two subsets (highly memorable images and low memorable images) following the same way in Sec. 4.2. In comparison, we also use our ICMscore predictor, proposed in Sec. 4.1, to divide the ICMD into three subsets LM, MM, and HM. The three subsets are then used to train image recognition models and continual learning models during replays.
Fig. A5 shows image recognition performance across the LM-model, the MM-model, and the HM-model. See Sec. 4.2 for the result discussion. In Fig. A5, we present the continual learning performance of various models with LM, MM, and HM images for replays. See Sec. 4.3 for the result discussion.
((a))
((b))
Figure A4: Performance comparison in Top-1 Accuracy between image classification models trained on HM, MM, and LM images using different memorability predictors. (a) shows results using our ICMscore predictor from Sec. 4.1, while (b) presents results from the regressor proposed in [5]. Within each subplot, from the top to the bottom, top-1 accuracy is reported for models trained on LM images (blue), HM images (green), and MM images (yellow). The chance is 20% for the 5-way classification (dash lines produced by the random model in red). The error bar is the standard error calculated over 5 independent runs with different random seeds.
((c))
((d))
Figure A5: Continual learning performance of AI models using HM, MM, and LM images for replays (a) shows results using our ICMscore predictor from Sec. 4.1, while (b) presents results from the regressor proposed in [5]. Within each subplot, from left to right, the performance of continual learning models using HM (green), MM (yellow), and LM (blue) images for replays is presented in Acc (left), BWT (middle), and Forgetting (right). See Sec. 4.3 for the introduction to the evaluation metrics and the continual learning models. The error bars are the standard errors calculated over 5 independent runs with different random seeds.
A.5 Performance of Models Using ICMscore without Time Penalty
We modify ICMscore by removing the time penalty, defining $s_{i}^{\prime}=1$ (correct) and $s_{i}^{\prime}=0$ (incorrect), while keeping ICMscore(I)$=\frac{\sum_{i=1}^{n}s_{i}}{n}$. Following the dataset splitting protocol from Sec. 4.2, we train image recognition and continual learning models on LM, MM, and HM subsets.
Fig. A6(a) shows that removing the time penalty reduces the top-1 recognition accuracy by 2% compared to our best LM model introduced in Sec. 4.2. Similarly, the best continual learning performance in Fig. A6(b) degrades by 3% compared to our LM continual learning model introduced in Sec. 4.3. Overall, removing the time penalty leads to narrowing the performance gap among LM, MM and HM models in both object recognition and continual learning tasks. Notably, the key finding remains consistent: the LM subset helps train a better recognition model while mitigating forgetting.
Figure A6: Performance of Models Using ICMscore without time penalty. (a) Top-1 accuracy comparison for image classification models trained on LM (blue), HM (green), and MM (yellow) images. The dashed line indicates the 20% chance level. (b) Continual learning performance showing Accuracy (left), BWT (middle), and Forgetting (right) for models using HM (green), MM (yellow), and LM (blue) replays. Error bars represent standard error over 5 independent runs. Evaluation metrics are detailed in Sec. 4.3.